# Discrete Distribution Networks

### v2

@@ 4.2 Ablation Study
In Table 2, we demonstrate the impact of different numbers of output nodes ($K$) and various techniques on the network. Interestingly, despite the substantial difference between having and not having the Split-and-Prune technique in the toy example, as shown in Figure 5, the performance in the ablation study without Split-and-Prune is not as poor as one might expect. We hypothesize that this is due to the Chain Dropout forcing all dead nodes to receive gradient guidance, preventing the network from generating poor results that are unoptimized. A particular case is when $K=2$, where the representational space of BinDDN is already sufficiently compact. The use of Chain Dropout in this case tends to result in more blurred generated images. Therefore, we did not employ Chain Dropout when $K=2$.
@@ Appendix A Experimental Results on Generative Performance
Conditional DDN for Image-to-Image Tasks. In the realm of image-to-image tasks, we extend our Discrete Distribution Networks (DDN) to a conditional setting, resulting in the conditional DDN. In this architecture, the condition is directly fed into the network during training in the beginning of each stage. The condition serves as an informative guide, significantly reducing the model’s generative space and, consequently, the complexity of the modeling task.
Those conditions inherently carry rich information, enabling the network to produce higher-quality samples. Through this conditional design, the DDN can leverage the abundant information contained in the conditions to generate more accurate and detailed images as shown in fig. 10.
Figure 10: Conditional DDN performing coloring and edge-to-RGB tasks. Benefiting from the reduction of the generative space by the condition, DDN is capable of generating high-quality images of $256\times 256$ resolution. Columns 4 and 5 display the generated results under the guidance of other images, where the produced image strives to adhere to the style of the guided image as closely as possible while ensuring compliance with the condition.
@@
Demonstration of Generation and Reconstruction Quality. figs. 13, 14 and 15 illustrate the generation and reconstruction results in various ablation experiments using the DDN model.
@@ Appendix B Further Demonstrations on Zero-Shot Conditional Generation
Utilizing CLIP as Conditioning Guidance. As illustrated in fig. 16, we leverage CLIP Radford et al. (2021) along with corresponding prompts as conditioning cues. We use a Discrete Distribution Network (DDN) model, which is exclusively trained on the FFHQ dataset, to yield Zero-Shot Conditional Generation (ZSCG) results. Remarkably, DDN can generate corresponding images under the guidance of CLIP without the necessity for gradient computations and only requiring a single inference pass.
ZSCG with Multiple Conditions. In fig. 17, we illustrate the operation of ZSCG under the combined action of two Guided Samplers: Inpainting and CLIP. Each sampler operates under its own specific condition. The inpainting sampler utilizes a mask to cover the areas where the CLIP prompt acts. Specifically, “wearing sunglasses” masks the eyes, “wearing a hat” masks the upper half of the face, and “happy person” masks the lower half. For each Discrete Distribution Layer (DDL), both samplers assign a rank to the generated images, corresponding to the degree of match to their respective conditions– the better the match, the higher the rank. We assign a weight to each sampler, which affects the assigned ranking. In this case, both samplers have a weight of 0.5. To promote diversity in the generated samples, we randomly select one image from the top two ranked by their condition-matching scores, serving as the output for that DDL layer.
@@ Appendix C Latent Analysis
Semantic Performance of Latents. We explored the semantic capabilities of DDN latents through a classification experiment on the MNIST dataset. Given the inherent tree structure of DDN’s latents, we employed a decision tree classification method, using fine-tuning data to assign class votes to nodes in the tree. For unassigned nodes, their class is inherited from the closest ancestor node with a class. We fine-tuned DDN’s latent decision tree using various numbers of labeled training set data, and the results on the test set are shown in table 3. All experiments in this table were conducted with Recurrence Iteration Paradigm’s UNet, which has approximately 407k parameters. However, computational complexity increases proportionally with the number of output levels (L). These experiments substantiate that DDN’s latents encompass meaningful semantic information.
@@ Appendix D More detailed experimental explanation
In fig. 19, we have expanded the “Illustration of the Split-and-Prune operation” by providing a schematic when $K$ increases to 15. This demonstrates that having a larger sample space results in an approximation that is closer to the target distribution.
@@
Figure 19: Illustration of the Split-and-Prune operation. For example in (a), the light blue bell-shaped curve represents a one-dimensional target distribution. The 5 “” under the x-axis are the initial values from a uniform distribution of 5 output nodes, which divide the entire space into 5 parts using midpoints between adjacent nodes as boundaries (i.e., vertical gray dashed lines). Each part corresponds to the range represented by this output node on the continuous space $x$. Below each node are three values: $P$ stands for the relative frequency of the ground truth falling within this node’s range during training; $Q$ refers to the probability mass of this sample (node) in the discrete distribution output by the model during the generation phase, which is generally equal for each sample, i.e., $1/K$. The bottom-most value denotes the difference between $P$ and $Q$. Colorful horizontal line segments represent the average probability density of $P$, $Q$ within corresponding intervals. In (b), the Split operation selects the node with the highest $P$ (circled in red). In (c), the Prune operation selects the node with the smallest $P$ (circled in red). In (d), through the combined effects of loss and Split-and-Prune operations, the distribution of output nodes moves towards final optimization. From the observed results, the KL divergence ($KL(P||Q)$) consistently decreases as the operation progresses, and the yellow line increasingly approximates the light blue target distribution. Finally, we show the distributions of the initial and final stages when the number of output nodes $K=15$ in (e) and (f). Due to the increased representational space (higher resolution), in (f), the probability density distribution $Q$ (yellow line segment) is closer to the lightblue target distribution than in the case of $K=5$.
Figure 20: Toy examples for two-dimensional data generation. The numerical values at the bottom of each figure represent the Kullback-Leibler (KL) divergence. Due to phenomena such as “dead nodes” and “density shift”, the application of Gradient Descent alone fails to properly fit the Ground Truth (GT) density. By employing the Split-and-Prune strategy, the density map looks the same as Real Samples. In the experiment, we use $K=10,000$ discrete nodes to emulate the probability distribution of GT density. Each node encompasses two parameters, $x$ and $y$, initialized from a uniform distribution. Each experiment consists of 100,000 iterations, where in each iteration an L2 loss is calculated based only on the node closest to the GT. The GT density map is converted into a discrete distribution with bins of size $100\times 100$, which is then used to calculate the KL divergence against the discrete distribution represented by these nodes. The KL divergence of Split-and-Prune is even lower than that of the Real Samples. This is because our algorithm has been exposed to 100,000 sets of GT data, thus it better reflects the GT distribution compared to the ‘Real Samples’, which are drawn only 10,000 times from the GT distribution.
Appendix E Limitations
@@
  • Difficulty with complex data: While random choice can generate decent facial data, DDNs struggle with complex data, such as ImageNet Krizhevsky et al. (2012). One possible solution is to learn an additional prior model dedicated to sampling output nodes.
  • computational burden of zero-shot conditional generation: ZSCG requires $L\times K$ forward passes through the guided model, where $L$ is the number of layers and $K$ is the number of possible outputs per layer. When the guided model itself is computationally expensive, this results in significant computational overhead and prolonged generation time. However, since the discrimination process is parallelizable across the $K$-dimensional output space, batching techniques can be employed to mitigate latency. In addition, further research will be conducted to reduce the number of guided model calls during the ZSCG process.

### v3

@@ 4.2 Ablation Study
In Table 2, we demonstrate the impact of different numbers of output nodes ($K$) and various techniques on the network. Interestingly, despite the substantial difference between having and not having the Split-and-Prune technique in the toy example, as shown in fig. 17, the performance in the ablation study without Split-and-Prune is not as poor as one might expect. We hypothesize that this is due to the Chain Dropout forcing all dead nodes to receive gradient guidance, preventing the network from generating poor results that are unoptimized. A particular case is when $K=2$, where the representational space of Taiji-DDN is already sufficiently compact. The use of Chain Dropout in this case tends to result in more blurred generated images. Therefore, we did not employ Chain Dropout when $K=2$.
@@ 5 Conclusion
Figure 6: Hierarchical Generation Visualization of DDN. We trained a DDN with output level $L=3$ and output nodes $K=8$ per level on MNIST dataset, its latent hierarchical structure is visualized as recursive grids. The large image in the middle is the average of all the generated images. Each sample with a colored border represents an intermediate generation image: a blue border indicates generation by the first layer, and a green border by the second layer. The samples within the surrounding grid of each colored-bordered sample are refined versions generated conditionally based on it (enclosed by the same color frontier). The small samples without colored borders are the final generated images. More detailed visualization of $L=4$ is presented in the appendix fig. 19.
@@ 6 Acknowledgements
The work was supported by National Science and Technology Major Project of China (2023ZD0121300).
@@ Appendix A Experimental Results on Generative Performance
Conditional DDN for Image-to-Image Tasks. In the domain of image-to-image tasks, we extend our Discrete Distribution Networks (DDN) to a conditional setting, resulting in the conditional DDN. The image condition is fed directly into the network during training at the beginning of each stage. The condition serves as an informative guide, significantly reducing the task’s generative space and, consequently, mitigating the modeling task’s complexity. Through this conditional design, the DDN can leverage the abundant information contained in the conditions to generate more accurate and detailed images as shown in fig. 8.
Figure 8: Conditional DDN performing coloring and edge-to-RGB tasks. Benefiting from the reduction of the generative space by the condition, DDN is capable of generating high-quality images of $256\times 256$ resolution. Columns 4 and 5 display combination of conditional generation and ZSCG, the generated results under the guidance of other images, where the produced image strives to adhere to the style of the guided image as closely as possible while ensuring compliance with the condition.
@@
Demonstration of Generation and Reconstruction Quality. Figures 12, 13 and 14 illustrate the results of generation and reconstruction in various ablation experiments using the DDN model.
@@
Implementation Details of Split-and-Prune. In algorithm 1, the Split and Prune operations are executed simultaneously. During implementation, the newly split nodes occupy the positions of the pruned nodes in the tensors, ensuring that the total number of network parameters remains constant and the GPU memory usage is fixed. It is important to note that the algorithm must not only operate on the output parameters but also apply the same operations to the corresponding state variables in the optimizer. This ensures consistency between the parameters being optimized and their associated state information throughout the training process.
Comparison with other novel generative models. We conduct a qualitative comparison with a recent work, the Idempotent Generative Network (IGN) Shocher et al. (2024), accepted by ICLR 2024. Since IGN was only experimented on the CelebA dataset and did not release its code, our comparison is limited on the CelebA dataset. As depicted in fig. 11, our DDN demonstrate better generation capabilities over IGN.
(a) DDN (Ours), FID=35.4
(b) Idempotent Generative Network (IGN), FID=39
@@ Appendix B Further Demonstrations on Zero-Shot Conditional Generation
Utilizing CLIP as Conditioning Guidance. As illustrated in fig. 15, we use CLIP Radford et al. (2021) along with the corresponding prompts as conditions. We use a DDN model, which is exclusively trained on the FFHQ dataset, to yield Zero-Shot Conditional Generation (ZSCG) results. The results indicate that, DDN can generate corresponding images under the guidance of CLIP without the necessity for gradient computations.
ZSCG with Multiple Conditions. In fig. 16, we illustrate the operation of ZSCG under the combined action of two Guided Samplers: Inpainting and CLIP. Each sampler operates under its own specific condition. The inpainting sampler utilizes a mask to cover the areas where the CLIP prompt acts. Specifically, “wearing sunglasses” masks the eyes, “wearing a hat” masks the upper half of the face, and “happy person” masks the lower half.
For each Discrete Distribution Layer (DDL), both samplers assign a rank to every generated image, corresponding to the degree of match to their respective conditions– the better the match, the higher the rank. We apply a weight to each sampler, which represents their influence on the final assigned ranking. In this case, both samplers have a weight of 0.5. To promote diversity in the generated samples, we randomly select one image from the top two ranked generated images, serving as the output for that DDL layer.
@@ Appendix C Latent Analysis
Semantic Performance of Latents. We explored the semantic capabilities of DDN latents through a classification experiment on the MNIST dataset. Given the inherent tree structure of DDN’s latents, we employed a decision tree classification method, using fine-tuning data to assign class votes to nodes in the tree. For unassigned nodes, their class is inherited from the closest ancestor node with an assigned class. We fine-tuned DDN’s latent decision tree using various numbers of labeled training set data, and the results on the test set are shown in table 3. All experiments in this table were conducted with Recurrence Iteration Paradigm’s UNet, which has approximately 407k parameters. These experiments substantiate that DDN’s latents encompass meaningful semantic information.
@@
Figure 17: Toy examples for two-dimensional data generation. The numerical values at the bottom of each figure represent the Kullback-Leibler divergence, $D_{KL}$. Due to phenomena such as “dead nodes” and “density shift”, applying Gradient Descent alone fails to properly fit the Ground Truth (GT) density. By employing the Split-and-Prune strategy, the density map looks the same as Real Samples. In the experiment, we use $K=10,000$ discrete nodes to emulate the probability distribution of the GT density. Each node encompasses two parameters, $x$ and $y$, initialized from a uniform distribution. Each experiment consists of $00010\times K=100,$ 000 iterations, wherein each iteration, an L2 loss is calculated based only on the node closest to the GT. To compute the KL divergence, the GT density map is converted into a discrete distribution with bins of size $100\times 100$, which is then used to calculate the $D_{KL}$ against the discrete distribution represented by these nodes. Notably, the KL divergence of Split-and-Prune is even lower than that of the Real Samples. This is because our algorithm has been exposed to $10\times K$ samples of the GT distribution, thus it better reflects the GT distribution compared to the Real Samples, which are drawn only $K$ times from the GT distribution.
(a) Famous quote and its translation about Taiji
(b) Tree structure of Taiji
Figure 18: The Taiji-DDN exhibits a surprising similarity to the ancient Chinese philosophy of Taiji. Records of Taiji can be traced back to the I Ching (Book of Changes) from the late 9th century BC, often described by the quote on the left (a) that explains the universe’s generation and transformation. This description coincidentally also summarizes the generation process and the transformations in the generative space of Taiji-DDN. Moreover, the diagram (b) from the book Tom (2013) bears a closely resemblance to the tree structure of DDN’s latent fig. 1(b). Therefore, we have named the DDN with $K=2$ as Taiji-DDN.
@@ Appendix D More detailed experimental explanation
In fig. 20, we have expanded the “Illustration of the principle behind the Split-and-Prune operation” by providing a schematic when $K$ increases to 15. This demonstrates that a larger sample space yields an approximation closer to the target distribution.
@@
Figure 20: Illustration of the principle behind the Split-and-Prune operation. For example in (a), the light blue bell-shaped curve represents a one-dimensional target distribution. The 5 “” under the x-axis are the initial values from a uniform distribution of 5 output nodes, which divide the entire space into 5 parts using midpoints between adjacent nodes as boundaries (i.e., vertical gray dashed lines). Each part corresponds to the range represented by this output node on the continuous space $x$. Below each node are three values: $P$ stands for the relative frequency of the ground truth falling within this node’s range during training; $Q$ refers to the probability mass of this sample (node) in the discrete distribution output by the model during the generation phase, which is generally equal for each sample, i.e., $1/K$. The bottom-most value denotes the difference between $P$ and $Q$. Colorful horizontal line segments represent the average probability density of $P$, $Q$ within corresponding intervals. In (b), the Split operation selects the node with the highest $P$ (circled in red). In (c), the Prune operation selects the node with the smallest $P$ (circled in red). In (d), through the combined effects of loss and Split-and-Prune operations, the distribution of output nodes moves towards final optimization. From the observed results, the KL divergence ($KL(P||Q)$) consistently decreases as the operation progresses, and the yellow line increasingly approximates the light blue target distribution. Finally, we show the distributions of the initial and final stages when the number of output nodes $K=15$ in (e) and (f). Due to the increased representational space, the generated probability distribution $Q$, represented by the yellow line segment in (f), is closer to the light blue target distribution than in the case of $K=5$.
Appendix E Limitations And Future Work
@@
  • The $K^{L}$ output space is insufficiently large to represent complex distributions.: We are currently developing a new approach to enhance the efficiency of high-dimensional data representation by expanding the output space from $K^{L}$. This involves dividing an image into $N$ patches, independently selecting the optimal patch from $K$ candidates for each patch. These selected patches are then merged into a complete image, which serves as the output of the current DDL layer and is input into the next layer. Consequently, this increases the output space to $K^{N\cdot L}$.
  • The Split-and-Prune algorithm continuously discards trained parameters.: Discarding parameters that have undergone extensive training is not a wise choice, particularly when scaling up the model. The goal of the Split-and-Prune algorithm is to balance the frequency at which each node is sampled during training, similar to maintaining load balance among experts in Mixture-of-Experts (MoE) models Liu et al. (2024). A potential solution to address this issue is the Loss-Free Balancing method Wang et al. (2024). After computing the L2 distance between each node’s output and the ground truth, Loss-Free Balancing first applies a node-wise bias to these distances. By dynamically updating the bias of each node based on its recent load, this method ensures that nodes with lower sampling frequencies receive a lower bias, thereby increasing their chances of being sampled. Consequently, Loss-Free Balancing helps maintain a balanced distribution of node sampling without discarding parameters.
@@
  • Computational burden of zero-shot conditional generation: ZSCG requires $L\times K$ forward passes through the guided model, where $L$ is the number of layers and $K$ is the number of possible outputs per layer. When the guided model itself is computationally expensive, this results in significant computational overhead and prolonged generation time. However, since the discrimination process is parallelizable across the $K$-dimensional output space, batching techniques can be employed to mitigate latency. In addition, it is not necessary for every layer to be guided by the condition signal, further research will be conducted to reduce the number of guided model calls during the ZSCG process.